OpenStudy (anonymous):

if I am trying to find the x and y intercept, how would I with "y=x-5"

jimthompson5910 (jim_thompson5910):

to find the y-intercept, replace x with 0 and evaluate what do you get?

OpenStudy (anonymous):

y=-5

jimthompson5910 (jim_thompson5910):

the y-intercept is (0,-5) which means there is a point at -5 on the y axis where the graph goes through 

jimthompson5910 (jim_thompson5910):

now plug in y = 0 and solve for x. This will give you the x-intercept

OpenStudy (anonymous):

what formula do I use?

jimthompson5910 (jim_thompson5910):

y = x-5 0 = x-5 ... replace y with 0 now isolate x. To do so, you add 5 to both sides (to undo the subtraction of 5)

OpenStudy (anonymous):

so 5 = x?

jimthompson5910 (jim_thompson5910):

yep x = 5 which means the x-intercept is (5,0) basically 5 on the x axis